# Probing quark-lepton correlation in GUTs with high-precision neutrino measurements

### v1

@@ Abstract
GUTs unify quarks and leptons into same representations and predict correlations between their masses and mixing. We take new data of JUNO and perform numerical scans to explore the flavor space compatible with data in SO(10) GUTs. The quark-lepton correlation shows the preference of normal ordering for light neutrino masses, predicts favored region of the CP-violating phase in neutrino oscillations, and classifies GUT models based their testability in neutrinoless double beta decay experiments. The quark-lepton correlation predicts mass spectrum of right-handed neutrinos, pointing to the energy scale of baryon and lepton number violation and providing sources for baryogenesis. We emphasize that, as the high precision measurements of neutrino physics is coming, the quark-lepton correlation will provide increasingly important role in the testability of GUTs, complementary to the proton decay measurement.
@@ I INTRODUCTION
Grand Unified Theories (GUTs) do not just unify three gauge interactions of the Standard Model (SM), but also unify Yukawa interactions of quarks and leptons. Of particular interests are the $SO(10)$ GUTs Fritzsch:1974nn where all SM matter fields are arranged in $16$-plets of $SO(10)$ and, as a consequence, all fermion masses and mixing are correlated Chanowitz:1977ye ; Ramond:1979py ; Harvey:1981hk . While proton decays provides a smoking gun signature of GUTs Weinberg:1979sa ; Wilczek:1979hc ; Weinberg:1980bf ; Weinberg:1981wj ; Sakai:1981pk ; Dimopoulos:1981dw ; Ellis:1981tv , the quark-lepton correlation of masses and mixing might verify the unification in the Yukawa sector. Furthermore, as right-handed neutrinos (RHNs) are embedded into $16$-plets, their mass spectrum and Yukawa couplings are predicted by the quark-lepton correlation, and successful leptogenesis become testable in the seesaw framework.
@@
The historical intertwining between neutrino measurements and GUTs dates back to the early 1980’s, when the first generation experiments represented by IMB and KamiokaNDE were operated (see in review Langacker:1980js ). Current best limits on proton decay is set by the second generation singularly comprised of Super-Kamiokande Abe:2014mwa ; Miura:2016krn . The latter is most famous for its observation of atmospheric neutrino oscillation in 1998 Super-Kamiokande:1998kpq .
@@
JUNO An:2015jdp , Hyper-K Acciarri:2015uup and DUNE Abe:2018uyc are three main large-scale neutrino oscillation measurements in this stage. They will continue on proton decay measurements to test GUTs by pushing the sensitivities of one order of magnitude. In the meantime, the most important goals are the determination of the neutrino mass hierarchy (normal $m_{1}<m_{2}<m_{3}$ or inverted $m_{3}<m_{1}<m_{2}$) and measurement of the Dirac-type CP-violating phase $\delta$. The octant of $\theta_{23}$ (1st octant $0<\theta_{23}<45^{\circ}$) and (2nd octant $45^{\circ}<\theta_{23}<90^{\circ}$) will be confirmed and measurements of other oscillation parameters with precision up to sub-percent level will be performed.
@@
JUNO released the first data on the measurement of $\sin^{2}\theta_{12}$ and $\Delta m^{2}_{21}$ after two months of running, with precision improved by a factor of 1.6 relative to the combination of all previous measurements JUNO:2025gmd . In the stage of high-precision neutrino data, the quark-lepton correlation provides tool of increasing significance to test GUTs complementary to proton decay. We will use the quark-lepton correlation in SO(10) GUTs as a criteria to explore the preferred distribution of unresolved physical observables, including neutrino mass ordering, the favored range of the Dirac CP-violating phase, and effective mass in neutrinoless double beta decay. These quantities will all be tested in the precision era of neutrino physics. In addition, the quark-lepton correlation predicts right-handed neutrino mass spectrum, which touches to BLB-L breaking scale and provides sources for baryon-antibaryon asymmetry in the observed Universe. The rest of this paper is organized as follows. Section II gives an overview on Yukawa couplings in SO(10) GUTs. Numerical analysis based on quark-lepton correlation is performed in Section III. We summarise and discuss our results in Section IV.
@@ II OVERVIEW ON YUKAWAS
Here, ${10}_{H}$, ${120}_{H}$ can be either real or complex, which will be denoted with subscripts $\mathbb{R}$ and $\mathbb{C}$, respectively, and $\overline{126}_{H}$ is always complex following their representation properties. Complexifying the Higgs induces couplings with the conjugates $({10}_{H}^{\mathbb{C}})^{*}$ and $({120}_{H}^{\mathbb{C}})^{*}$. They can be forbidden by imposing the Peccei-Quinn (PQ) symmetry $U(1)_{PQ}$ Babu:1992ia ; Bajc:2005zf , which gives a compelling explanation of the strong CP problem Peccei:1977hh . The PQ symmetry is a global $U(1)$ symmetry in which fermions and Higgses transform as
@@
Following the symmetry breaking from GUTs to the SM, these Higgses are eventually decomposed to eight electroweak doublets in total,
@@
While the general Yukawa couplings in Eq. (1) leaves plenty room to fit data of fermion masses and mixing, economical models have been considered to reduce the number of free parameters. We discuss the economical and realistic Higgs content to be included in this work. First, one single Higgs multiplet (either ${10}_{H}$, $\overline{126}_{H}$ or ${120}_{H}$) is not enough since the prediction of generate fermion masses and no flavor mixing in either quark and lepton sector. To explain tiny masses for light neutrinos, $\overline{126}_{H}$ must be included since it is the only Higgs involving a SM singlet with charge BL=2B-L=2, which is necessary to generate Majorana masses for neutrinos via seesaw mechanisms. The minimal extension is by adding a real $10^{\mathbb{R}}_{H}$ into the theory. However, it cannot explain the large hierarchy between $m_{b}$ and $m_{t}$ Bajc:2005zf . On the other hand, the model with $(\overline{126}_{H},120_{H})$ is found to be unable to fit all the fermion masses and mixing angles in the three generation case Joshipura:2011nn ; Dueck:2013gca . Then, we are left with a few realistic choices in fitting data, either the minimal Higgs content $({10}_{H}^{\mathbb{C}},\overline{126}_{H}^{\mathbb{C}})$ or a complete list $({10}_{H},\overline{126}_{H},{120}_{H})$. Among them, we will focus on three compelling models.
@@
      This one introduces the minimal Higgs contents among the realistic models Babu:1992ia . Imposing $U(1)_{PQ}$ forbid terms with $({10}_{H}^{\mathbb{C}})^{*}$ and thus only $Y_{10}$ and $Y_{126}$ are present Mummidi:2021anm ; Ohlsson:2019sja ; Babu:2015bna ; Altarelli:2013aqa . Without loss of generality, we rotate the flavor basis to keep $Y_{126}$ always real and diagonal. There are $6\times 2+3$ real parameters in $Y_{10}$ and $Y_{126}$. Additional free parameters includes the relative size and phases among $c^{u,d}_{10,126}$. In total, we have 19 free parameters in the Yukawa sector.
  • M2 – $({10}_{H}^{\mathbb{R}},\overline{126}{}_{H}^{\mathbb{C}},{120}_{H}^{\mathbb{R}})$ with SCPV.
@@
  • M3 – $({10}_{H}^{\mathbb{C}},\overline{126}{}_{H}^{\mathbb{C}},{120}_{H}^{\mathbb{C}})$ with $U(1)_{\rm PQ}\times Z_{2}^{P}$.
@@ III NUMERICAL ANALYSIS
It is worth noting that NuFIT with SK atmospheric data prefers $\theta_{23}$ in the first octant ($\theta_{23}<45^{\circ}$) and that without SK atmospheric data prefers the second octant ($\theta_{23}>45^{\circ}$). We have taken both data into account to check the preference of the $\theta_{23}$ octant in those GUT models. The up-to-date experimental constraint on the Dirac CP-violating phase $\delta$ is still weak, and thus is not included in the $\chi^{2}$. Instead, we treat $\delta$ as a prediction of the model, together with other predictions including the effective masses appearing in $\beta$ decay $m_{\beta}$, neutrinoless double-$\beta$ decay $m_{\beta\beta}$, neutrino mass sum contributing in cosmology $\sum m_{i}$, and right-handed neutrino mass spectrum $M_{1},M_{2},M_{3}$.
We first comment on the preference of neutrino mass ordering in SO(10) GUTs. We have done scan in both normal ordering (NO) and inverted ordering (IO) with the same computation power. In MI with the IO, M1 cannot minimizes $\chi^{2}/n_{\rm obs}$ below 280; In M2 and M3, we have found points marginally consistent with data with $\chi^{2}/n_{\rm obs}=35$ and $\chi^{2}/n_{\rm obs}=19$ respectively in the IO. In general, scanning for IO is more difficult than the NO case. In all three models, NO is highly favored by the quark-lepton correlation.
In the NO, general results are shown in Fig. 1, with benchmarks presented in Appendix B. We summarize the results below.
  • By taking data of $\theta_{23}$ in both octants, we find a slight preference on the 1st octant in M2, but no significant preference of the octant in M1 and M3.
  • Distributions of $\delta$ are different in three models. $\delta$ is restricted in the region $90^{\circ},+90^{\circ}]$ in M2, preferred in the same region in M3, and nearly flatly distributed in the whole parameter space in M1.
  • $m_{1}$ is predicted around in the region $[0.002,0.01]$ eV in M1, $m_{1}\lesssim 0.001$ eV in M2 and is mainly distributed in $10^{-4},10^{-2}]$ eV in M3.
  • $m_{\beta\beta}$ in M1 is predicted in a narrow regime $[0.004,0.01]$ eV, which can be tested in the future $0\nu 2\beta$ experiments. Values of $m_{\beta\beta}$ in M1 is smaller, in the region $[0.002,0.004]$ eV, below the experimental sensitivity. The predicted region in M3 overlaps with both models.
@@
We discuss the predictions of RHN mass spectrum and potentially consequent phenomenon. These models are predicted very different mass spectrums for RHNs. We have confirmed the large hierarchy of RHN masses in M2 which has been studied in Babu:2024ahk , where the lightest one can be as small as $10^{4}$ GeV and the heaviest one can reach $10^{15}$ GeV. However, more points show the lightest mass $M_{1}$ located in the interval $10^{6},3\cdot 10^{7}]$ GeV. This regime has not been discovered before. Both M1 and M3 shows much smaller hierarchies. M1 gives a clear prediction on $M_{1}$ around $10^{9},10^{10}]$ GeV and the highest one $M_{3}\sim[10^{12},10^{13}]$ GeV. where most points give $M_{1}/M_{3}\sim 10^{-3}$. Predictions in M3 are more diverse, although a lot of them gives $M_{1}\sim 10^{10}$ GeV and $M_{3}\sim 10^{13}$ GeV. This result is consistent with former results Fu:2022lrn , but we have found more parameter space.
@@
We further comment that a clear prediction of RHN mass spectrum serves as the basis for numerous interesting research avenues. We briefly address them below. RHNs provide an explanation of baryon-antibaryon asymmetry via the mechanism of leptogenesis. This is quantitatively calculated via the quark-lepton correlation in SO(10) GUTs. A successful thermal leptogenesis requires $M_{1}$ heavier than $10^{9}$ GeV Davidson:2002qv , which applies for most points in M1 and M3. However, this condition is satisfied in M2, and inspired alternative scenario, e.g., $N_{2}$ leptogenesis DiBari:2010ux might succeed in some parameter space Babu:2025wop . In a perturbative theory, all RHN masses should not be heavier than the BLB-L energy scale. That means this scale should be heavier than $M_{3}\sim 10^{13}$ GeV for most points in M1 and M3. In M2, the heaviest RHN can be as heavy as $10^{15}$ GeV. A BLB-L scale above this value requires special RG effects to be included in the gauge unification.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (NSFC) under Grant Nos. 12205064, 12347103, and Zhejiang Provincial Natural Science Foundation of China under Grant No. LDQ24A050002.
Refer to caption Figure 1: Predictions of GUT models M1, M2 and M3.
@@ Appendix A Details of fitting procedure
Regarding VEVs of Higgs doublets in ${10}_{H},\overline{126}_{H},{120}_{H}$, without loss of generality, $v_{126}^{u}$ can be taken to be real, and renmaining VEVs are generally complex. Moreover, SM singlet S VEV $v_{S}$ can also be real, concretely we write
@@
where these coefficients were obtained from VEVs
$\{\langle h_{10}^{u}\rangle,\langle h_{10}^{d}\rangle,\langle h_{126}^{u}\rangle,\langle h_{126}^{d}\rangle,$
$\langle h_{120_{1}}^{u}\rangle,\langle h_{120_{15}}^{u}\rangle,\langle h_{120_{1}}^{d}\rangle,\langle h_{120_{15}}^{d}\rangle\}$
$=$ $\{c_{10}^{u},c_{10}^{d},c_{126}^{u},c_{126}^{d},$
$c_{120_{1}}^{u},c_{120_{1}}^{d},c_{120_{15}}^{u},c_{120_{15}}^{d}\}\times v_{\rm EW}\,.$ (9)
Furthermore, we define the following Yukawa matrices and real parameters
@@
With the above preparation, three models in this paper can be reparameterized, which are shown below.
@@
    Due to CP invariance, $Y_{10},Y_{126},Y_{120}$ are all real matrices, equivalently $S,D,A$ are real matrices and $M_{R}$ is a real symmetric matrix. Here, we work in a basis where $Y_{126}$ is diagonal and positive, resulting $D$ is diagonal and positive. $S$ is real symmetric and $A$ is real antisymmetric. $r_{1},r_{4},v_{R}$ are real parameters and $\alpha,\lambda,\gamma,\omega$ are phases.
@@
    Here, we follow the convention in Ref. Fu:2022lrn , as CP invariance above the GUT scale and spontaneous CP violation (SCPV) below the GUT scale, $S,D,A$ are real matrices, resulting $Y_{u},Y_{d},Y_{e},Y_{\nu}$ are Hermitian matrices and $M_{R}$ is a real symmetric matrix. Consistent with previous models, we still choose a basis where $D$ is diagonal and positive, $S$ is real symmetric and $A$ is real antisymmetric. $r_{1},r_{2},r_{3},r_{4},r_{5},v_{R}$ are real parameters.
@@
Data (best-fits $\pm 1\sigma$) for quark Yukawas and mixing parameters, as well as charged lepton Yukawas, are listed below, taken from Babu:2016bmy with RG effect running up to $M_{\rm GUT}=2\times 10^{16}$ GeV.
@@
For charged lepton yukawas, we have quoted error to be 1% uncertainty in their masses to relax the difficulty induced by the tiny experimental errors of charged lepton masses Babu:2024ahk . Oscillation data except $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$ in the normal ordering are taken from NuFIT 6.0 Esteban:2024eli . In the normal ordering,
@@
$\Delta m^{2}_{31}$ $=(2.534\pm 0.024)\times 10^{-3}{\rm eV}^{2}\,$, (16)
for IC19 without SK atmospheric data and
$\Delta m^{2}_{31}$ $=(2.513\pm 0.020)\times 10^{-3}{\rm eV}^{2}\,$, (17)
for IC24 with SK atmospheric data. In the inverted ordering, those without or with SK atmospheric data are respectively given by
@@
$\Delta m^{2}_{32}$ $=(-2.510\pm 0.025)\times 10^{-3}{\rm eV}^{2}\,$, (18)
$\sin^{2}\theta_{23}$ $=0.550\pm 0.014\,,\;\sin^{2}\theta_{13}=0.02231\pm 0.00057$
$\Delta m^{2}_{32}$ $=(-2.484\pm 0.020)\times 10^{-3}{\rm eV}^{2}\,.$ (19)
NuFIT data of $\sin^{2}\theta_{12}$ and $\Delta m^{2}_{21}$ in the inverted ordering keep the same,
@@
The scan procedure is performed as below.

### v2

@@ Abstract
GUTs unify quarks and leptons into same representations and predict correlations between their masses and mixing. We perform numerical scans in SO(10) GUTs to explore the flavor space with new data of JUNO taken into account. The quark-lepton correlation shows the preference of normal ordering for light neutrino masses, predicts favored region of the CP-violating phase in neutrino oscillations, and classifies GUT models based on their testability in neutrinoless double beta decay experiments. The quark-lepton correlation predicts mass spectrum of right-handed neutrinos, pointing to the energy scale of baryon and lepton number violation and providing sources for baryogenesis. We emphasize that, as high precision measurements of neutrino physics are coming, the quark-lepton correlation will provide increasingly important role in the testability of GUTs, complementary to proton decay measurements.
@@ I INTRODUCTION
Grand Unified Theories (GUTs) do not just unify three gauge interactions of the Standard Model (SM), but also unify Yukawa interactions of quarks and leptons. Of particular interest are the $SO(10)$ GUTs Fritzsch:1974nn where all SM matter fields are arranged in $16$-plets of $SO(10)$ and, as a consequence, all fermion masses and mixing are correlated Chanowitz:1977ye ; Ramond:1979py ; Harvey:1981hk . While the proton decay provides a smoking gun signature of GUTs Weinberg:1979sa ; Wilczek:1979hc ; Weinberg:1980bf ; Weinberg:1981wj ; Sakai:1981pk ; Dimopoulos:1981dw ; Ellis:1981tv , the quark-lepton correlation of masses and mixing might verify the unification of matters via the Yukawa sector. Furthermore, as right-handed neutrinos (RHNs) are embedded into $16$-plets, their mass spectrum and Yukawa couplings are predicted by the quark-lepton correlation, and successful leptogenesis becomes testable in the seesaw framework.
@@
The historical intertwining between neutrino measurements and GUTs dates back to the early 1980’s, when the first generation experiments represented by IMB and KamiokaNDE were operated (see in review Langacker:1980js ). Current best limits on proton decay are set by the second generation singularly comprised of Super-Kamiokande Abe:2014mwa ; Miura:2016krn . The latter is most famous for its observation of atmospheric neutrino oscillation in 1998 Super-Kamiokande:1998kpq .
@@
JUNO An:2015jdp , Hyper-K Acciarri:2015uup and DUNE Abe:2018uyc are three main large-scale neutrino oscillation measurements in this stage. They will continue on proton decay measurements to test GUTs and push sensitivities by one order of magnitude. In the meantime, the most important goals are the determination of the neutrino mass ordering (normal $m_{1}<m_{2}<m_{3}$ or inverted $m_{3}<m_{1}<m_{2}$) and measurement of the Dirac-type CP-violating phase $\delta$. The octant of $\theta_{23}$ (1st octant $0<\theta_{23}<45^{\circ}$) and (2nd octant $45^{\circ}<\theta_{23}<90^{\circ}$) will be confirmed and measurements of other oscillation parameters with precision up to sub-percent level will be carried out.
@@
JUNO released the first data on the measurement of $\sin^{2}\theta_{12}$ and $\Delta m^{2}_{21}$ after two months of running, with precision improved by a factor of 1.6 relative to the combination of all previous measurements JUNO:2025gmd . In the stage of high-precision neutrino data, the quark-lepton correlation provides a tool of increasing significance to test GUTs complementary to proton decay. We will use the quark-lepton correlation in $SO(10)$ GUTs as a criterion to explore the preferred distribution of unresolved physical observables, including neutrino mass ordering, the favored range of the Dirac CP-violating phase, and effective mass in neutrinoless double-beta ($0\nu\beta\beta$) decay. These quantities will all be tested in the precision era of neutrino physics. In addition, the quark-lepton correlation predicts right-handed neutrino mass spectrum, which touches to the BLB-L breaking scale and provides sources for baryon-antibaryon asymmetry in the observed Universe. The rest of this paper is organized as follows. Section II gives an overview on Yukawa couplings in $SO(10)$ GUTs. Numerical analysis based on quark-lepton correlation is performed in Section III. We summarize and discuss our results in Section IV.
@@ II OVERVIEW ON YUKAWAS
Here, ${10}_{H}$, ${120}_{H}$ can be either real or complex, which will be denoted with superscripts $\mathbb{R}$ and $\mathbb{C}$, respectively, and $\overline{126}_{H}$ is always complex following its representation property. Complexifying the Higgs induces couplings with the conjugates $({10}_{H}^{\mathbb{C}})^{*}$ and $({120}_{H}^{\mathbb{C}})^{*}$. They can be forbidden by imposing the global Peccei-Quinn (PQ) symmetry $U(1)_{\rm PQ}$ Babu:1992ia ; Bajc:2005zf , which gives a compelling explanation of the strong CP problem Peccei:1977hh . Fermions and Higgses under $U(1)_{\rm PQ}$ transform as
@@
Following the symmetry breaking from GUTs to the SM, these Higgs fields are eventually decomposed to eight electroweak doublets in total,
@@
While the general Yukawa couplings in Eq. (1) leaves plenty of room to fit data of fermion masses and mixing, economical models have been considered to reduce the number of free parameters. We discuss the economical and realistic Higgs content to be included in this work. First, one single Higgs multiplet (either ${10}_{H}$, $\overline{126}_{H}$ or ${120}_{H}$) is not enough, since it fails to generate fermion masses and no flavor mixing in either quark or lepton sector. To explain tiny masses for light neutrinos, $\overline{126}_{H}$ must be included since it is the only Higgs involving a SM singlet with charge BL=2B-L=2, which is necessary to generate Majorana masses for neutrinos via seesaw mechanism. The minimal extension is by adding a real $10^{\mathbb{R}}_{H}$ into the theory. However, it cannot explain the large hierarchy between $m_{b}$ and $m_{t}$ Bajc:2005zf . On the other hand, the model with $(\overline{126}_{H},120_{H})$ is found to be unable to fit all the fermion masses and mixing angles in the three generation case Joshipura:2011nn ; Dueck:2013gca . Then, we are left with a few realistic choices for fitting data, either the minimal Higgs content $({10}_{H}^{\mathbb{C}},\overline{126}_{H}^{\mathbb{C}})$ or a complete list $({10}_{H},\overline{126}_{H},{120}_{H})$. Among them, we will focus on three compelling models.
@@
      This one introduces the minimal Higgs contents among the realistic models Babu:1992ia . Imposing $U(1)_{\rm PQ}$ forbids terms with $({10}_{H}^{\mathbb{C}})^{*}$ and thus only $Y_{10}$ and $Y_{126}$ are present Mummidi:2021anm ; Ohlsson:2019sja ; Babu:2015bna ; Altarelli:2013aqa . Without loss of generality, we rotate the flavor basis to keep $Y_{126}$ always real and diagonal. There are $6\times 2+3$ real parameters in $Y_{10}$ and $Y_{126}$. Additional free parameters includes the relative size and phases among $c^{u,d}_{10,126}$. In total, we have 19 free parameters in the Yukawa sector.
[…]
@@
[…]
@@ III NUMERICAL ANALYSIS
NuFIT provides four versions of global data, which we list in Appendix A and denote as N1, N2, I1 and I2, depending on which mass ordering is assumed and whether SK atmospheric data is included. Namely, N1 and N2 represent global fits with and without SK atmospheric data in the normal ordering, and I1 and I2 are those in the inverted ordering. The up-to-date experimental constraint on the Dirac CP-violating phase $\delta$ is still weak, and thus is not included in the $\chi^{2}$. Instead, we treat $\delta$ as a prediction of the model, together with other predictions including the lightest neutrino mass $m_{\text{lightest}}$, the effective masses appearing in $\beta$ decay $m_{\beta}$, $0\nu\beta\beta$ decay $m_{\beta\beta}$, and right-handed neutrino masses $M_{1},M_{2},M_{3}$ (from lightest to heaviest).
We perform numerical scan for all three models M1, M2, M3 with four versions of data respectively accounted. Same procedures of scan are performed in all cases. All viable points for $\chi^{2}/n_{\rm obs}<10$ in these models are presented in Appendix A, in Figs. 3 and 4 for normal and inverted ordering, respectively. Representative benchmarks in these models are presented in Appendix B.
We first report results on the prediction of neutrino mass ordering. In model M1, no viable points in the inverted ordering are found to fit data in our scan. In models M2 and M3, although both normal and inverted ordering can fit data, the former case is highly favored due to its efficiency in the scan. Scan efficiencies of models in fitting both mass ordering are compared in Fig. 1.
Figure 1: Viable points (p) per core hour (ch) obtained in M1, M2 and M3 with $\chi^{2}/n_{\rm obs}<10$ in each scan.
Below we will mainly focus on the normal mass ordering. Note that data N1 and N2 prefer $\theta_{23}$ in the 1st octant ($\theta_{23}<45^{\circ}$) and 2nd octant ($\theta_{23}>45^{\circ}$), respectively. Thus, we can further check the preference of the $\theta_{23}$ octant in models. Given points in Fig. 3, we count points in certain intervals normalized by the corresponding weight $e^{-\chi^{2}/2}$ and obtain the preferred regimes of observables in models. General results on predictions of $\theta_{23}$, $\delta$ and RHN masses are shown in Fig. 2 and summarized as follows. By taking data of $\theta_{23}$ in both octants, we find a slight preference on the 1st octant in M2, but no significant preference of the octant in M1 and M3. Distributions of $\delta$ are different in three models. $\delta$ is mainly distributed in the region $70^{\circ},+70^{\circ}]$ in M2, extended to $90^{\circ},+90^{\circ}]$ in M3 and spanned into most of the parameter space in M1.
@@
We discuss the predictions of RHN mass spectrum and potentially consequent phenomenon. These models predict very different mass spectrums for RHNs. We have confirmed the large hierarchy of RHN masses in M2 which has been studied in Babu:2024ahk , where the lightest one can be as small as $10^{4}$ GeV and the heaviest one can reach $10^{15}$ GeV. However, more points show the lightest mass $M_{1}$ span in the interval $4\cdot 10^{4},3\cdot 10^{7}]$ GeV. This regime has not been discovered before. Both M1 and M3 show much smaller hierarchies. M1 gives a clear prediction on $M_{1}$ around $10^{9},10^{10}]$ GeV and the highest one $M_{3}\sim[10^{12},10^{13}]$ GeV, where most points give $M_{1}/M_{3}\sim 10^{-3}$. Predictions in M3 are more diverse, although a lot of them gives $M_{1}\sim 10^{10}$ GeV and $M_{3}\sim 10^{13}$ GeV. This result is consistent with former results Fu:2022lrn , but we have found more parameter space.
@@
A clear prediction of RHN mass spectrum serves as the basis for numerous interesting research avenues. We briefly address them below. RHNs provide an explanation of baryon-antibaryon asymmetry via the mechanism of leptogenesis. This is quantitatively calculated via the quark-lepton correlation in $SO(10)$ GUTs. A successful thermal leptogenesis requires $M_{1}$ heavier than $10^{9}$ GeV Davidson:2002qv , which applies for most points in M1 and M3. However, this condition is not satisfied in M2, and inspires alternative scenarios, e.g., $N_{2}$ leptogenesis DiBari:2010ux , might succeed in some parameter space Babu:2025wop . In a perturbative theory, all RHN masses should not be heavier than the BLB-L energy scale. That means this scale should be heavier than $M_{3}\sim 10^{13}$ GeV for most points in M1 and M3. In M2, the heaviest RHN can be as heavy as $10^{15}$ GeV. A BLB-L scale above this value requires special RG effects to be included in the gauge unification.
Figure 2: Predicted distributions of some observables of models M1, M2 and M3 in the normal order in fitting data N1 and N2 in the normal ordering. RHN masses $M_{1}$, $M_{2}$ and $M_{3}$ are marked in red, green and blue, respectively in the bottom panel.
@@ Acknowledgements.
This work was supported by Zhejiang Provincial Natural Science Foundation of China (No. LDQ24A050002) and National Natural Science Foundation of China (NSFC) (Nos. 12535007 and 12547104).
@@ Appendix A Details of fitting procedure
After GUT symmetry breaking, ${10}_{H},\overline{126}_{H},{120}_{H}$ are decomposed to a series of electroweak doublets of the SM gauge symmetry, denoted in Eq. (3). They mix and the lightest massless one matches to the SM Higgs. After the SM Higgs gains the VEV $v$, triggering the electroweak symmetry breaking, the SM fermions obtain masses. We parameterize VEVs of electroweak doublets as
$\{\langle h_{10}^{u}\rangle,\langle h_{10}^{d}\rangle,\langle h_{120_{1}}^{u}\rangle,\langle h_{120_{1}}^{d}\rangle\}$
$=$ $\{c_{10}^{u},c_{10}^{d},c_{120_{1}}^{u},c_{120_{1}}^{d}\}\times\frac{v}{\sqrt{2}}\,;$
$\{\langle h_{126}^{u}\rangle,\langle h_{126}^{d}\rangle,\langle h_{120_{15}}^{u}\rangle,\langle h_{120_{15}}^{d}\rangle\}$
$=$ $\{c_{126}^{u},c_{126}^{d},c_{120_{15}}^{u},c_{120_{15}}^{d}\}\times\sqrt{\frac{3}{2}}v\,$, (8)
where all $c_{\rm rep}^{u,d}$ are dimensionless and complex parameters, and overall factors in front of $v$ are induced by representation decomposition in group theory, which are irrelevant to our fit. We further parameterize $c_{\rm rep}^{u,d}$ by removing unphysical parameters. Without loss of generality, $c_{126}^{u}$ can be rotated to be real, and remaining coefficients are generally complex. Concretely we write
@@
Moreover, the SM singlet scalar VEV $v_{S}$ can also taken to be real. Therefore, Yukawa matrices and parameters are written as
@@
With the above preparation, three models in this paper can be reparameterized, as below.
@@
    Due to the CP symmetry, $Y_{10},Y_{126},Y_{120}$ are all real, equivalently $S,D,A$ are real matrices and $M_{R}$ is a real symmetric matrix. Here, we work in a basis where $D$ is diagonal and positive. $S$ is real symmetric and $A$ is real antisymmetric. $r_{1},r_{4},v_{R}$ are real parameters and $\alpha,\lambda,\gamma,\omega$ are phases.
@@
    We follow the convention in Ref. Fu:2022lrn , as CP invariance above the GUT scale and spontaneous CP violation (SCPV) below the GUT scale, $S,D,A$ are real matrices, resulting $Y_{u},Y_{d},Y_{e},Y_{\nu}$ are Hermitian matrices and $M_{R}$ is a real symmetric matrix. Consistent with previous models, we still choose a basis where $D$ is diagonal and positive, $S$ is real symmetric and $A$ is real antisymmetric. $r_{1},r_{2},r_{3},r_{4},r_{5},v_{R}$ are real parameters.
@@
Table 3: Parameters counting in three models.
Label Free parameters
$\begin{array}[]{c}{\tt M1}\\ (n_{\rm para}=19)\end{array}$ $\begin{array}[]{c}\{r_{1},r_{2}\}\in(-10,10),\ m_{0}\in(0,1)\ {\rm eV},\ \omega\in[0,2\pi),\\ D:\{D_{11},D_{22},D_{33}\}\in(0,1),\\ S:\{S_{11}e^{is_{1}},S_{12}e^{is_{2}},S_{13}e^{is_{3}},S_{22}e^{is_{4}},S_{23}e^{is_{5}},S_{33}e^{is_{6}}\}\in(-10,10)e^{i[0,2\pi)}\end{array}$
$\begin{array}[]{c}{\tt M2}\\ (n_{\rm para}=19)\end{array}$ $\begin{array}[]{c}\{r_{1},r_{4}\}\in(-10,10),~m_{0}\in(0,1)\ {\rm eV},~\{\alpha,\lambda,\gamma,\omega\}\in[0,2\pi),\\ D:\{D_{11},D_{22},D_{33}\}\in(0,1),\\ S:\{S_{11},S_{12},S_{13},S_{22},S_{23},S_{33}\}\in(-10,10),\\ A:\{A_{12},A_{13},A_{23}\}\in(-10,10)\end{array}$
$\begin{array}[]{c}{\tt M3}\\ (n_{\rm para}=18)\end{array}$ $\begin{array}[]{c}\{r_{1},r_{2},r_{3},r_{4},r_{5}\}\in(-10,10),~m_{0}\in(0,1)\ {\rm eV},\\ D:\{D_{11},D_{22},D_{33}\}\in(0,1),\\ S:\{S_{11},S_{12},S_{13},S_{22},S_{23},S_{33}\}\in(-10,10),\\ A:\{A_{12},A_{13},A_{23}\}\in(-10,10)\end{array}$
Data ($\text{bf}\pm 1\sigma$) for quark Yukawas and mixing parameters, as well as charged lepton Yukawas, are listed below, taken from Babu:2016bmy with RG effect running up to $M_{\rm GUT}=2\times 10^{16}$ GeV.
@@
For charged lepton Yukawas, we have quoted error to be 1% uncertainty in their masses to relax the difficulty induced by the tiny experimental errors of charged lepton masses Babu:2024ahk . Oscillation data except $\Delta m^{2}_{21}$ and $\sin^{2}\theta_{12}$ in the normal ordering are taken from NuFIT 6.0 Esteban:2024eli . We apply following four versions of data: normal ordering for IC24 with SK atmospheric data (denoted as N1),
$\Delta m^{2}_{31}$ $=(2.513\pm 0.020)\times 10^{-3}{\rm eV}^{2}\,;$ (16)
normal ordering for IC19 without SK atmospheric data (N2),
@@
$\Delta m^{2}_{31}$ $=(2.534\pm 0.024)\times 10^{-3}{\rm eV}^{2}\,;$ (17)
inverted ordering with SK atmospheric data (I1),
$\sin^{2}\theta_{23}$ $=0.550\pm 0.014\,,\;\sin^{2}\theta_{13}=0.02231\pm 0.00056$
$\Delta m^{2}_{32}$ $=(-2.484\pm 0.020)\times 10^{-3}{\rm eV}^{2}\,;$ (18)
and inverted ordering without SK atmospheric data (I2),
@@
$\Delta m^{2}_{32}$ $=(-2.510\pm 0.025)\times 10^{-3}{\rm eV}^{2}\,.$ (19)
In the inverted ordering, both I1 and I2 support $\theta_{23}$ in the 2nd octant. NuFIT data of $\sin^{2}\theta_{12}$ and $\Delta m^{2}_{21}$ in inverted ordering keep the same,
@@
In all of our scans, the same procedure is performed as below.